# Gauge and Scalar Boson Mediated Proton Decay in a Predictive SU(5) GUT Model

### v1

@@ Abstract
We assess proton decay signatures in the simplest viable $SU(5)$ model with regard to constraints on parameters governing the Standard Model fermion mass spectrum. Experimental signals for all eight two-body proton decay processes result from exchange of two gauge bosons, a single scalar leptoquark, or their combination. Consequently, it enables us to delve into an in-depth anatomy of proton decay modes and anticipate future signatures. Our findings dictate that observing a proton decay into $p\to\pi^{0}e^{+}$ indicates gauge boson mediation, with the potential for observation of $p\to\eta^{0}e^{+}$ mode. Alternatively, if decay is through $p\to K^{+}\overline{\nu}$ process, it is mediated by a scalar leptoquark, possibly allowing the observation of $p\to\pi^{0}\mu^{+}$. Detection of both $p\to\pi^{0}e^{+}$ and $p\to K^{+}\overline{\nu}$ could enhance $p\to\pi^{0}\mu^{+}$ through constructive interference. The model predicts inaccessibility of $p\to\pi^{+}\overline{\nu}$, $p\to\eta^{0}\mu^{+}$, $p\to K^{0}e^{+}$, and $p\to K^{0}\mu^{+}$, regardless of the dominant mediation type, in the coming decades. In summary, through a comprehensive analysis of proton decay signals, gauge coupling unification, and fermion masses and mixing, we precisely constrain the parameter space of the $SU(5)$ model in question.
@@ 2 Model
The full scalar potential reads
@@
Our one-loop level gauge coupling unification analysis reveals that fields $\Phi_{10}$ and $\Phi_{1}$ need to be degenerate in mass and heavy, i.e., of the order of $10^{11}$ GeV, whereas $\Phi_{3}$ and $\Phi_{6}$ also need to be degenerate in mass but light, i.e., of the order of $10$ TeV, if $M_{\mathrm{GUT}}$ is to be at, or exceed, existing experimental limits on partial proton decay lifetimes. Moreover, under the assumption that parameters $M_{\Sigma_{6}}$, $M_{\Sigma_{3}}$, and $M_{\Sigma_{1}}$ are all positive, the masses of $\Sigma_{6}$, $\Sigma_{3}$, and $\Sigma_{1}$ fields are degenerate and, as such, are not constrained by the gauge coupling unification requirement, at all. It turns out, however, that phenomenological viability of neutrino masses requires $M_{\Sigma_{1}}$ to be of the order of $M_{\Phi_{1}}$, as we demonstrate later on.
We implement one-loop level gauge coupling unification analysis in order to find the largest possible value of unification scale $M_{\mathrm{GUT}}$ and associated value of $\alpha_{\mathrm{GUT}}=g^{2}_{\mathrm{GUT}}/(4\pi)$ for fixed values of $M_{\Phi_{1}}$ and $M_{\Sigma_{1}}$. Again, since $M_{\Phi_{1}}\sim M_{\Phi_{10}}$ and $M_{\Sigma_{1}}\sim M_{\Sigma_{3}}\sim M_{\Sigma_{6}}$, the only remaining parameters we need to vary within our unification analysis are the masses of $\Phi_{3},\Phi_{6}\in 35_{H}$, $\phi_{1},\phi_{8}\in 24_{H}$, and $\Lambda_{3}\in 5_{H}$, while the relevant SM input parameters are $M_{Z}=91.1876$ GeV, $\alpha_{S}(M_{Z})=0.1193\pm 0.0016$, $\alpha^{-1}(M_{Z})=127.906\pm 0.019$, and $\sin^{2}\theta_{W}(M_{Z})=0.23126\pm 0.00005$ Agashe:2014kda . Since we also require $M_{\Lambda_{3}}$ to be at, or above, $3\times 10^{11}$ GeV due to the scalar mediated proton decay requirements Dorsner:2012uz , the only masses one actually needs to vary in our approach are $M_{\Phi_{3}}$, $M_{\Phi_{6}}$, $M_{\phi_{1}}$, and $M_{\phi_{8}}$.
@@
One thing to note is that the value of $M_{\mathrm{GUT}}$ is independent of $M_{\Sigma_{1}}$ due to degeneracy of fields in $15_{F}+\overline{15}_{F}$, in agreement with our previous discussion. Also, to produce $M_{\mathrm{GUT}}$ values, as given in Fig. 2, we need to have four multiplets to be exactly at $M$ scale. These multiplets are $\Phi_{3},\Phi_{6}\in 35_{H}$ and $\phi_{1},\phi_{8}\in 24_{H}$. If the mass of any of these four multiplets is simply shifted up, the unification scale $M_{\mathrm{GUT}}$ would go down. This means that we are discussing the most conservative scenario with regard to the proton decay signatures via gauge boson exchange. Also, one can treat scale $M$, for all practical purposes, as a geometric mean of masses $M_{\Phi_{3}}$, $M_{\Phi_{6}}$, $M_{\phi_{1}}$, and $M_{\phi_{8}}$ when one studies impact of its change on $M_{\mathrm{GUT}}$.
@@ 3.1 Charged fermion sector
$\mathcal{L}_{Y}\supset$ $\begin{pmatrix}u_{i}&\Sigma^{u}\end{pmatrix}\begin{pmatrix}\begin{array}[]{c|c}\sqrt{2}v_{5}(Y^{u}_{ij}+Y^{u}_{ji})&-\frac{5v_{24}}{4}Y^{c}_{i}\\ \hline\cr 0&M_{\Sigma_{3}}\end{array}\end{pmatrix}\begin{pmatrix}u^{C}_{j}\\ \overline{\Sigma}^{u}\end{pmatrix}$ (3.30)
$+\begin{pmatrix}d_{i}&\Sigma^{d}\end{pmatrix}\begin{pmatrix}\begin{array}[]{c|c}\frac{1}{2}v_{5}Y^{d}_{ij}&-\frac{5v_{24}}{4}Y^{c}_{i}\\ \hline\cr\frac{1}{\sqrt{2}}v_{5}Y^{a}_{j}&M_{\Sigma_{3}}\end{array}\end{pmatrix}\begin{pmatrix}d^{C}_{j}\\ \overline{\Sigma}^{d}\end{pmatrix}$ (3.37)
@@ 4 Proton Decay
$p\rightarrow\pi^{+}\overline{\nu}$ $3.9\times 10^{33}$ years Super-Kamiokande:2013rwg -
@@ 5 Predictions
Again, Fig. 2 features contours of constant values for $M_{\mathrm{GUT}}/(10^{15}\,\mathrm{GeV})$ (red solid lines), $\alpha_{\mathrm{GUT}}$ (blue dot-dashed lines), and $m_{0}/(\sqrt{\Delta m^{2}_{31}}/2)$ (green solid lines) in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane for $M\geq 1$ TeV and $M\geq 10$ TeV scenarios, where $\log_{10}(M_{\Phi_{1}}/1\,\mathrm{GeV})\in[10.2,12.9]$ and $\log_{10}(M_{\Sigma_{1}}/1\,\mathrm{GeV})\in[8.4,14.2]$. The lattice spacing in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane for our analysis, in both directions, is 0.1 in units of $\log_{10}(M_{\Phi_{1}}/1\,\mathrm{GeV})$ and $\log_{10}(M_{\Sigma_{1}}/1\,\mathrm{GeV})$, respectively. Again, at each lattice point in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane we evaluate mass spectrum that leads to maximal possible value of $M_{\mathrm{GUT}}$ and determine associated $\alpha_{\mathrm{GUT}}$, $A_{SL}$, and $A_{SR}$. The mass spectrum that generates max ($M_{\mathrm{GUT}}$) is always such that $M_{\Phi_{3}}=M_{\Phi_{6}}=M_{\phi_{1}}=M_{\phi_{8}}=M$. We furthermore perform running of the SM charged fermion parameters at each lattice point and use associated values to perform numerical fit of all fermion masses. The outcome of the fit are all unitary transformations that provide transition from flavor to mass-eigenstate basis of the SM fermions and associated Yukawa coupling constants. Parameters that our procedure cannot determine are phases $\eta_{1}$, $\eta_{2}$, $\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$, $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$, where only $\eta_{1}$ and $\eta_{2}$ are relevant for the study of proton decay signatures.
@@
Once these ten points in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane are known, we choose among them one point with large value of $\alpha_{\mathrm{GUT}}$ (Q$(10^{11.3}\,\mathrm{GeV},10^{9.5}\,\mathrm{GeV})$) and another one with small value of $\alpha_{\mathrm{GUT}}$ (R$(10^{11.0}\,\mathrm{GeV},10^{12.7}\,\mathrm{GeV})$) and evaluate at both of these points gauge boson mediated proton decay widths for all eight channels, where we also vary $\eta_{1}$ and $\eta_{2}$ phases to find maximal and minimal values for these decay widths. We furthermore evaluate scalar leptoquark mediated proton decay widths for all eight channels at both Q and R, where, once again, we vary $\eta_{1}$ and $\eta_{2}$ to capture phase dependence. To find decay widths associated with scalar leptoquark mediation, we fix the mass of $\Lambda_{3}$ to have a partial lifetime for $p\to K^{+}\overline{\nu}$ as close as possible to the current experimental limit. We take $M_{\Lambda_{3}}=2.7\times 10^{12}$ GeV for $M\geq$ 1 TeV scenario. Again, this choice ensures that the scalar mediated proton decay signatures will be observed with certainty at future proton decay experiments.
@@
Upper panels in Fig. 3 are for proton decay signatures for points Q and R, whereas lower panels correspond to predictions for points ${}^{\prime}$ and ${}^{\prime}$. Note that the uncertainties in proton decay widths associated with scalar leptoquark mediation are much larger than those associated with gauge boson mediation. Again, in both cases these uncertainties originate solely from variation of two phases $\eta_{1}$ and $\eta_{2}$. We quantify these flavor uncertainties in Table IV, where we evaluate $\Delta=(\max(\tau_{p})-\min(\tau_{p}))/\min(\tau_{p})$ in % at points Q, R, ${}^{\prime}$, and ${}^{\prime}$ for each process and for both types of mediation.
@@
Table IV: Phase uncertainty $\Delta=(\max(\tau_{p})-\min(\tau_{p}))/\min(\tau_{p})$ in % at points Q, R, ${}^{\prime}$, and ${}^{\prime}$ for all eight channels and for both types of proton decay mediation.
@@
We have already stipulated that variation of $\eta_{2}$ does not affect gauge mediated proton decay signatures at all. To that end, we present in Fig. 4 contours of constant decay width for $p\to\pi^{0}\mu^{+}$ for both the gauge (blue) and scalar leptoquark (red) mediation in $\eta_{1}$-$\eta_{2}$ plane for points Q (upper panels) and ${}^{\prime}$ (lower panels). It is clear from Fig. 4 that the impact of $\eta_{2}$ on gauge boson mediated proton decay signatures is marginal, at best.
Figure 4: Contours of constant $\Gamma(p\to\pi^{0}\mu^{+})$ for the gauge boson (blue) and scalar leptoquark (red) mediation for points Q and ${}^{\prime}$, as indicated.
@@
The very last points in the parameter space of $M\geq 1$ TeV and $M\geq 10$ TeV scenarios in Fig. 2 that could be, at least in principle, probed by proton decay experiments are points A and ${}^{\prime}$, respectively, with common coordinates $(M_{\Phi_{1}},M_{\Sigma_{1}})=(10^{12.8}\,\mathrm{GeV},10^{12.9}\,\mathrm{GeV})$. The model yields $\tau_{p}(p\to\pi^{0}e^{+})=1.1\times 10^{36}$ years and $\tau_{p}(p\to\pi^{0}e^{+})=1.5\times 10^{35}$ years at points A and ${}^{\prime}$, respectively. $M$ is thus inversely proportional to predicted lifetime $\tau_{p}(p\to\pi^{0}e^{+})$ at point $(M_{\Phi_{1}},M_{\Sigma_{1}})=(10^{12.8}\,\mathrm{GeV},10^{12.9}\,\mathrm{GeV})$. We can accordingly infer that the current experimental limit on $\tau_{p}(p\to\pi^{0}e^{+})$ implies that $M\leq 65$ TeV, where $M$ can be identified, for all practical purposes, with geometric mean of the masses of $\Phi_{3}$, $\Phi_{6}$, $\phi_{1}$, and $\phi_{8}$. Viable parameter space of our model is nicely bounded from one side by experimental limit on $p\to\pi^{0}e^{+}$ and from the other side by neutrino mass scale perturbativity requirement. The $M=65$ TeV scenario is where these two bounds meet.
We can also place an accurate limit on $M_{\Lambda_{3}}$ within $M\geq 1$ TeV and $M\geq 10$ TeV scenarios for any lattice point in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane with the data on $p\to K^{+}\overline{\nu}$. For example, we find that $M_{\Lambda_{3}}\geq 1.8\times 10^{12}$ GeV, $M_{\Lambda_{3}}\geq 1.9\times 10^{12}$ GeV, $M_{\Lambda_{3}}\geq 1.8\times 10^{12}$ GeV, and $M_{\Lambda_{3}}\geq 1.7\times 10^{12}$ GeV for points Q, R, ${}^{\prime}$, ${}^{\prime}$, respectively.
@@
($c$) We remark that, in principle, the fermionic submultiplet $\Sigma_{3}$ does not need to be degenerate in mass with $\Sigma_{6}$ and $\Sigma_{1}$ if one does not require $M_{\Sigma_{3}},M_{\Sigma_{6}},M_{\Sigma_{1}}>0$. This might be problematic for our study since, as shown in Ref. Antusch:2023mqe , a lower mass of $\Sigma_{3}$, by itself, can significantly increase the unification scale. One might then expect that taking its mass close to the TeV scale would drastically increase $M_{\mathrm{GUT}}$. This, however, does not happen in our model due to the presence of additional colored light states. For example, if one sets $\Sigma_{3}$ to be very light, the maximal unification scale $M_{\mathrm{GUT}}$ is only a factor of 1.5 larger than what we show in Fig. 2. Setting $\Sigma_{3}$ light, however, opens up another issue. To explain it, let us, for example, focus on the third-generation mass. We get, from Eq. (3.47), that $m_{b}\sim m_{\tau}+\Delta m$, where $\Delta m\sim(5/4)(174\,\mathrm{GeV})(Y^{c}Y^{a})(v_{24}/M_{\Sigma_{3}})$. To reproduce correct bottom mass, using the data from Table II, one requires $\Delta m\sim 1$ GeV. Assuming that $\Sigma_{3}$ state indeed resides at the TeV scale and $v_{24}\sim 2\times 10^{16}$ GeV, we obtain $Y^{c}\sim 10^{-17}$, where $Y^{a}\sim 1$ is used to maximize the neutrino mass scale. We present, in this manuscript, scenario with mass-degenerate $\Sigma_{1}$, $\Sigma_{6}$, and $\Sigma_{3}$ states, which dictates $Y^{c}$ coupling to be of the order of the usual electron Yukawa coupling and leave thorough discussion of the light $\Sigma_{3}$ scenario for potential future publication.

### v2

@@ Abstract
We assess proton decay signatures in the simplest viable $SU(5)$ model with regard to constraints on parameters governing the Standard Model fermion mass spectrum. Experimental signals for all eight two-body proton decay processes result from exchange of two gauge bosons, a single scalar leptoquark, or their combination. Consequently, it enables us to delve into an in-depth anatomy of proton decay modes and anticipate future signatures. Our findings dictate that observing a proton decay into $p\to\pi^{0}e^{+}$ indicates gauge boson mediation, with the potential for observation of $p\to\eta^{0}e^{+}$ mode. Alternatively, if decay is through $p\to K^{+}\overline{\nu}$ process, it is mediated by a scalar leptoquark, possibly allowing the observation of $p\to\pi^{0}\mu^{+}$. Detection of both $p\to\pi^{0}e^{+}$ and $p\to K^{+}\overline{\nu}$ could enhance $p\to\pi^{0}\mu^{+}$ through constructive interference. The model predicts inaccessibility of $p\to\pi^{+}\overline{\nu}$, $p\to\eta^{0}\mu^{+}$, $p\to K^{0}e^{+}$, and $p\to K^{0}\mu^{+}$, regardless of the dominant mediation type, in the coming decades. In summary, through a comprehensive analysis of proton decay signals, gauge coupling unification, and fermion masses and mixing, we both accurately and precisely constrain the parameter space of the $SU(5)$ model in question.
@@ 2 Model
The scalar potential of the model reads
@@
Our one-loop level gauge coupling unification analysis reveals that fields $\Phi_{10}$ and $\Phi_{1}$ need to be degenerate in mass and heavy, i.e., of the order of $10^{11}$ GeV, whereas $\Phi_{3}$ and $\Phi_{6}$ also need to be degenerate in mass but light, i.e., of the order of $10$ TeV or less, if $M_{\mathrm{GUT}}$ is to be at, or exceed, existing experimental limits on partial proton decay lifetimes. The mass spectrum of fields in $\Phi=35_{H}$ that yields the highest possible value of $M_{\mathrm{GUT}}$ is thus $M_{\Phi_{3}}=M_{\Phi_{6}}\ll M_{\Phi_{10}}=M_{\Phi_{1}}$.
Also, under the assumption that parameters $M_{\Sigma_{6}}$, $M_{\Sigma_{3}}$, and $M_{\Sigma_{1}}$ are all positive, our gauge coupling unification study shows that fields $\Sigma_{6}$, $\Sigma_{3}$, and $\Sigma_{1}$ want to be almost perfectly mass-degenerate. This happens for the following two reasons. Firstly, Eq. (2.8) shows that $M_{\Sigma_{3}}$ is an arithmetic mean of $M_{\Sigma_{1}}$ and $M_{\Sigma_{6}}$. Secondly, the only field that has the correct $\beta$-function coefficients to increase $M_{\mathrm{GUT}}$ is $\Sigma_{3}$, whereas $\Sigma_{1}$ and $\Sigma_{6}$ can only decrease it, if they reside below $M_{\mathrm{GUT}}$. So, there are only two possible non-degenerate scenarios in agreement with Eq. (2.8) that one needs to consider to fully understand this issue. One scenario is when $M_{\Sigma_{1}}\ll M_{\Sigma_{3}}\sim M_{\Sigma_{6}}$ and the other is when $M_{\Sigma_{6}}\ll M_{\Sigma_{3}}\sim M_{\Sigma_{1}}$. It is clear that in both instances the GUT scale would be smaller that in the case when $M_{\Sigma_{6}}\sim M_{\Sigma_{3}}\sim M_{\Sigma_{1}}$ since the lightest state, i.e., either $\Sigma_{1}$ or $\Sigma_{6}$, would reduce the maximal allowed value of the GUT scale, whereas the other two states cannot subsequently affect $M_{\mathrm{GUT}}$ since they would enter the gauge coupling running at practically the same threshold, at some higher scale, and would thus, together with the lightest state, form a full $SU(5)$ multiplet. A preferred scenario that yields the largest $M_{\mathrm{GUT}}$ value and generates unification of gauge coupling constants turns out to be a scenario when $M_{\Sigma_{6}}$ is at most a factor of 10 below $M_{\Sigma_{3}}$, while $M_{\Sigma_{3}}$ is a factor $1/2$ below $M_{\Sigma_{1}}$, all in accordance with Eq. (2.8).
The fact that $\Sigma_{6}$, $\Sigma_{3}$, and $\Sigma_{1}$ want to be almost perfectly mass-degenerate implies that their common mass scale, in that regime, is not constrained by the gauge coupling unification requirement, at all. It turns out, however, that phenomenological viability of neutrino masses requires $M_{\Sigma_{1}}$ to be of the order of $M_{\Phi_{1}}$, as we demonstrate later on.
We implement one-loop level gauge coupling unification analysis in order to find the largest possible value of unification scale $M_{\mathrm{GUT}}$ and associated value of $\alpha_{\mathrm{GUT}}=g^{2}_{\mathrm{GUT}}/(4\pi)$ for fixed values of $M_{\Phi_{1}}$ and $M_{\Sigma_{1}}$. Again, since $M_{\Phi_{1}}=M_{\Phi_{10}}$ and $M_{\Sigma_{1}}\sim M_{\Sigma_{3}}\sim M_{\Sigma_{6}}$, the only parameters we need to vary within our unification analysis are the masses of $\Phi_{3},\Phi_{6}\in 35_{H}$, $\phi_{1},\phi_{8}\in 24_{H}$, and $\Lambda_{3}\in 5_{H}$, while the relevant SM input parameters are $M_{Z}=91.1876$ GeV, $\alpha_{S}(M_{Z})=0.1193\pm 0.0016$, $\alpha^{-1}(M_{Z})=127.906\pm 0.019$, and $\sin^{2}\theta_{W}(M_{Z})=0.23126\pm 0.00005$ Agashe:2014kda . Since we also require $M_{\Lambda_{3}}$ to be at, or above, $3\times 10^{11}$ GeV due to the scalar mediated proton decay requirements Dorsner:2012uz , the only masses one actually needs to vary in our approach are $M_{\Phi_{3}}$, $M_{\Phi_{6}}$, $M_{\phi_{1}}$, and $M_{\phi_{8}}$.
@@
One thing to note is that the value of $M_{\mathrm{GUT}}$ is independent of $M_{\Sigma_{1}}$ due to almost perfect mass-degeneracy of fields in $15_{F}+\overline{15}_{F}$, in agreement with our previous discussion. Also, to produce $M_{\mathrm{GUT}}$ values, as given in Fig. 2, we need to have four multiplets to be exactly at $M$ scale. These multiplets are $\Phi_{3},\Phi_{6}\in 35_{H}$ and $\phi_{1},\phi_{8}\in 24_{H}$. If the mass of any of these four multiplets is simply shifted up, the unification scale $M_{\mathrm{GUT}}$ would go down. This means that we are discussing the most conservative scenario with regard to the proton decay signatures via gauge boson exchange. Also, one can treat scale $M$, for all practical purposes, as a geometric mean of masses $M_{\Phi_{3}}$, $M_{\Phi_{6}}$, $M_{\phi_{1}}$, and $M_{\phi_{8}}$ when one studies impact of its change on $M_{\mathrm{GUT}}$.
@@ 3.1 Charged fermion sector
$\mathcal{L}_{Y}\supset$ $\begin{pmatrix}u_{i}&\Sigma^{u}\end{pmatrix}\begin{pmatrix}\begin{array}[]{c|c}\sqrt{2}v_{5}(Y^{u}_{ij}+Y^{u}_{ji})&-\frac{5v_{24}}{4}Y^{c}_{i}\\ \hline\cr 0&M_{\Sigma_{3}}\end{array}\end{pmatrix}\begin{pmatrix}u^{C}_{j}\\ \overline{\Sigma}^{u}\end{pmatrix}$
$+\begin{pmatrix}d_{i}&\Sigma^{d}\end{pmatrix}\begin{pmatrix}\begin{array}[]{c|c}\frac{1}{2}v_{5}Y^{d}_{ij}&-\frac{5v_{24}}{4}Y^{c}_{i}\\ \hline\cr\frac{1}{\sqrt{2}}v_{5}Y^{a}_{j}&M_{\Sigma_{3}}\end{array}\end{pmatrix}\begin{pmatrix}d^{C}_{j}\\ \overline{\Sigma}^{d}\end{pmatrix}$
@@ 4 Proton Decay
$p\rightarrow\pi^{+}\overline{\nu}$ $3.9\times 10^{32}$ years Super-Kamiokande:2013rwg -
@@ 5 Predictions
Again, Fig. 2 features contours of constant values for $M_{\mathrm{GUT}}/(10^{15}\,\mathrm{GeV})$ (red solid lines), $\alpha_{\mathrm{GUT}}$ (blue dot-dashed lines), and $m_{0}/(\sqrt{\Delta m^{2}_{31}}/2)$ (green solid lines) in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane for $M\geq 1$ TeV and $M\geq 10$ TeV scenarios, where $\log_{10}(M_{\Phi_{1}}/1\,\mathrm{GeV})\in[10.2,12.9]$ and $\log_{10}(M_{\Sigma_{1}}/1\,\mathrm{GeV})\in[8.4,14.2]$. The lattice spacing in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane for our analysis, in both directions, is 0.1 in units of $\log_{10}(M_{\Phi_{1}}/1\,\mathrm{GeV})$ and $\log_{10}(M_{\Sigma_{1}}/1\,\mathrm{GeV})$, respectively. Again, at each lattice point in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane we evaluate mass spectrum that leads to maximal possible value of $M_{\mathrm{GUT}}$ and determine associated $\alpha_{\mathrm{GUT}}$, $A_{SL}$, and $A_{SR}$. The mass spectrum that generates $\max(M_{\mathrm{GUT}})$ is always such that $M_{\Phi_{3}}=M_{\Phi_{6}}=M_{\phi_{1}}=M_{\phi_{8}}=M$. We furthermore perform running of the SM charged fermion parameters at each lattice point and use associated values to perform numerical fit of all fermion masses. The outcome of the fit are all unitary transformations that provide transition from flavor to mass-eigenstate basis of the SM fermions and associated Yukawa coupling constants. Parameters that our procedure cannot determine are phases $\eta_{1}$, $\eta_{2}$, $\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$, $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$, where only $\eta_{1}$ and $\eta_{2}$ are relevant for the study of proton decay signatures.
@@
Once these ten points in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane are known, we choose among them one point with large value of $\alpha_{\mathrm{GUT}}$ (Q$(10^{11.3}\,\mathrm{GeV},10^{9.5}\,\mathrm{GeV})$) and another one with small value of $\alpha_{\mathrm{GUT}}$ (R$(10^{11.0}\,\mathrm{GeV},10^{12.7}\,\mathrm{GeV})$), where R is singled out because it has value of $\alpha_{\mathrm{GUT}}^{2}/M^{4}_{\mathrm{GUT}}$ that is numerically almost identical to $\alpha_{\mathrm{GUT}}^{2}/M^{4}_{\mathrm{GUT}}$ evaluated at point Q. We subsequently evaluate at both of these points gauge boson mediated proton decay widths for all eight channels, where we also vary $\eta_{1}$ and $\eta_{2}$ phases to find maximal and minimal values for these decay widths. We furthermore evaluate scalar leptoquark mediated proton decay widths for all eight channels at both Q and R, where, once again, we vary $\eta_{1}$ and $\eta_{2}$ to capture phase dependence. To find decay widths associated with scalar leptoquark mediation, we fix the mass of $\Lambda_{3}$ to have a partial lifetime for $p\to K^{+}\overline{\nu}$ as close as possible to the current experimental limit. We take $M_{\Lambda_{3}}=2.7\times 10^{12}$ GeV for $M\geq$ 1 TeV scenario. Again, this choice ensures that the scalar mediated proton decay signatures will be observed with certainty at future proton decay experiments.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Upper panels in Fig. 3 are for proton decay signatures for points Q and R, whereas lower panels correspond to predictions for points Q and R. Note that the uncertainties in proton decay widths associated with scalar leptoquark mediation are much larger than those associated with gauge boson mediation. Again, in both cases these uncertainties originate solely from variation of two phases $\eta_{1}$ and $\eta_{2}$. We quantify these flavor uncertainties in Table IV, where we evaluate $\Delta=(\max(\tau_{p})-\min(\tau_{p}))/\min(\tau_{p})$ in % at points Q, R, Q, and R for each process and for both types of mediation.
@@
Table IV: Phase uncertainty $\Delta=(\max(\tau_{p})-\min(\tau_{p}))/\min(\tau_{p})$ in % at points Q, R, Q, and R for all eight channels and for both types of proton decay mediation.
@@
We have already stipulated that variation of $\eta_{2}$ does not affect gauge mediated proton decay signatures at all. To that end, we present in Fig. 4 contours of constant decay width for $p\to\pi^{0}\mu^{+}$ for both the gauge (blue) and scalar leptoquark (red) mediation in $\eta_{1}$-$\eta_{2}$ plane for points Q (upper panels) and Q (lower panels). It is clear from Fig. 4 that the impact of $\eta_{2}$ on gauge boson mediated proton decay signatures is marginal, at best.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 4: Contours of constant $\Gamma(p\to\pi^{0}\mu^{+})$ for the gauge boson (blue) and scalar leptoquark (red) mediation for points Q and Q, as indicated.
@@
The very last points in the parameter space of $M\geq 1$ TeV and $M\geq 10$ TeV scenarios in Fig. 2 that could be, at least in principle, probed by proton decay experiments are points A and A, respectively, with common coordinates $(M_{\Phi_{1}},M_{\Sigma_{1}})=(10^{12.8}\,\mathrm{GeV},10^{12.9}\,\mathrm{GeV})$. The model yields $\tau_{p}(p\to\pi^{0}e^{+})=1.1\times 10^{36}$ years and $\tau_{p}(p\to\pi^{0}e^{+})=1.5\times 10^{35}$ years at points A and A, respectively. $M$ is thus inversely proportional to predicted lifetime $\tau_{p}(p\to\pi^{0}e^{+})$ at point $(M_{\Phi_{1}},M_{\Sigma_{1}})=(10^{12.8}\,\mathrm{GeV},10^{12.9}\,\mathrm{GeV})$. We can accordingly infer that the current experimental limit on $\tau_{p}(p\to\pi^{0}e^{+})$ implies that $M\leq 65$ TeV, where $M$ can be identified, for all practical purposes, with geometric mean of the masses of $\Phi_{3}$, $\Phi_{6}$, $\phi_{1}$, and $\phi_{8}$. Viable parameter space of our model is nicely bounded from one side by experimental limit on $p\to\pi^{0}e^{+}$ and from the other side by neutrino mass scale perturbativity requirement. The $M=65$ TeV scenario is where these two bounds meet.
We can also place an accurate limit on $M_{\Lambda_{3}}$ within $M\geq 1$ TeV and $M\geq 10$ TeV scenarios for any lattice point in $M_{\Phi_{1}}$-$M_{\Sigma_{1}}$ plane with the data on $p\to K^{+}\overline{\nu}$. For example, we find that $M_{\Lambda_{3}}\geq 1.8\times 10^{12}$ GeV, $M_{\Lambda_{3}}\geq 1.9\times 10^{12}$ GeV, $M_{\Lambda_{3}}\geq 1.8\times 10^{12}$ GeV, and $M_{\Lambda_{3}}\geq 1.7\times 10^{12}$ GeV for points Q, R, Q, R, respectively.
Our predictions, up to this point, have been based on the one-loop level gauge coupling unification analysis. We can, in principle, perform the same analysis at the two-loop level. To that end we present the two-loop level prediction for the unification scale at point A with coordinates $(M_{\Phi_{1}},M_{\Sigma_{1}})=(10^{12.8}\,\mathrm{GeV},10^{12.9}\,\mathrm{GeV})$ for the $M\geq 1$ TeV scenario presented in the upper panel of Fig. 2. The one-loop level analysis at point A yields $M_{\mathrm{GUT}}=1.97\times 10^{16}$ GeV and $\alpha_{\mathrm{GUT}}=0.0457$, whereas the two-loop level maximization of $M_{\mathrm{GUT}}$ at the same point yields $M_{\mathrm{GUT}}=5.94\times 10^{16}$ GeV and $\alpha_{\mathrm{GUT}}=0.0610$. The mass spectrum that generates this two-loop level unification scenario is $M_{\Phi_{3}}=M_{\Phi_{6}}=M_{\phi_{1}}=M_{\phi_{8}}=1$ TeV, $M_{\Sigma_{6}}=4.75\times 10^{11}$ GeV, $M_{\Delta_{3}}=10^{12}$ GeV, $M_{\Sigma_{3}}=3.97\times 10^{12}$ GeV, $M_{\Phi_{1}}=M_{\Phi_{10}}=10^{12.8}$ GeV, and $M_{\Sigma_{1}}=10^{12.9}$ GeV, where we use the two-loop $\beta$-function coefficients given in Appendix 3. We accordingly note that the ratio $\alpha_{\mathrm{GUT}}^{2}/M_{\mathrm{GUT}}^{4}$ at point A changes by a multiplicative factor of 46 when we go from the one-loop to two-loop level result. If we assume that the short-distance coefficients will be enhanced by a factor of 1.1 due to increase of the gauge coupling value in the case of the two-loop level analysis, we can see that the proton decay lifetimes will be scaled by, at most, a factor of 40 for point A. What is important, though, is that the pattern of proton decay signatures shown in Fig. 3 will not be changed if one goes from the one-loop to two-loop level analysis.
@@
($c$) We remark that, in principle, the fermionic submultiplet $\Sigma_{3}$ does not need to be almost perfectly degenerate in mass with $\Sigma_{6}$ and $\Sigma_{1}$ if one does not require $M_{\Sigma_{3}},M_{\Sigma_{6}},M_{\Sigma_{1}}>0$. This might be problematic for our study since, as shown in Ref. Antusch:2023mqe , a lower mass of $\Sigma_{3}$, by itself, can significantly increase the unification scale. One might then expect that taking its mass close to the TeV scale would drastically increase $M_{\mathrm{GUT}}$. This, however, does not happen in our model due to the presence of additional colored light states. For example, if one sets $\Sigma_{3}$ to be very light, the maximal unification scale $M_{\mathrm{GUT}}$ is only a factor of 1.5 larger than what we show in Fig. 2. Setting $\Sigma_{3}$ light, however, opens up another issue. To explain it, let us, for example, focus on the third-generation mass. We get, from Eq. (3.17), that $m_{b}\sim m_{\tau}+\Delta m$, where $\Delta m\sim(5/4)(174\,\mathrm{GeV})(Y^{c}Y^{a})(v_{24}/M_{\Sigma_{3}})$. To reproduce correct bottom mass, using the data from Table II, one requires $\Delta m\sim 1$ GeV. Assuming that $\Sigma_{3}$ state indeed resides at the TeV scale and $v_{24}\sim 2\times 10^{16}$ GeV, we obtain $Y^{c}\sim 10^{-17}$, where $Y^{a}\sim 1$ is used to maximize the neutrino mass scale. We present, in this manuscript, scenario with mass-degenerate $\Sigma_{1}$, $\Sigma_{6}$, and $\Sigma_{3}$ states, which dictates $Y^{c}$ coupling to be of the order of the usual electron Yukawa coupling and leave thorough discussion of the light $\Sigma_{3}$ scenario for potential future publication.